OpenStudy (anonymous):

simplify 2√5+3√5-√5 6√5 5√15 4√5 5√5

OpenStudy (anonymous):

Same system... \[(?+?-?)\sqrt{5}\]

OpenStudy (gabbyalicorn):

2√5+3√5−√5= 4√5

OpenStudy (anonymous):

thats the thing i dont get square roots

OpenStudy (anonymous):

thanks @gabbyalicorn

OpenStudy (anonymous):

They are just numbers, you can factor them out, like any numbers.

OpenStudy (anonymous):

10*5+5 = 5(10+1)

OpenStudy (gabbyalicorn):

Wait, you should just do the steps with @Redcan Cuz I feel the bad vibes....o-o

OpenStudy (anonymous):

lmao no bad vibes smh

OpenStudy (anonymous):

yea @Redcan i think i get where your coming from

OpenStudy (anonymous):

After you see the factoring, it makes all the question look the same.

OpenStudy (anonymous):

yea i get it now